# DiffSHEG: A Diffusion-Based Approach for Real-Time Speech-driven Holistic 3D Expression and Gesture Generation

### 09 Jan 2024 v1

@@ 4.3 Qualitative Results
Qualitative comparisons. We highly recommend readers watch our demo video to get an intuitive sense of qualitative results. Overall, in addition to better speech-motion alignment, we observe more realistic, agile and diverse expressions and gestures of ours than that of baselines on both datasets (Figure 4,  6,  7,  5). On the BEAT dataset, our method shows more agile gestures than CaMN. The gestures of baseline diffusion-based methods all suffer different extents of jittering. The jittering extents from high to low are DiffGesture, DiffuseStyleGesture, and LDA. Both our method and CaMN can generate smooth motions. However, CaMN exhibits much slower and less diverse motion while ours is as agile as the real motion. For expression, CaMN shows inaccurate lip sync while other methods all look fine. Interestingly, we find that the CaMN barely blinks, while our method can generate good expressions with a reasonable frequency of blinks (Figure 4). Other baselines show a more frequent blinking, corresponding to higher diversity scores. However, ours has the most similar blink frequency as the real data. On the SHOW dataset, our method shows more agile gestures than TalkSHOW and more diverse expressions and gestures (Figure 67). This might be due to the fact that Talkshow utilizes VQ-VAE to encode the gestures, which may filter the high-frequency gestures, leading to slow motion. Also, because VQ-VAE tokenizes the gestures into limited numbers, its gestures show lower diversity than ours. The LS3DCG suffers from jittering arms, which may be because their deterministic CNN structure has difficulty approximating the many-to-many mapping. Note that we retrain the TalkSHOW for qualitative comparison because the results of TalkSHOW and LS3DCG with author-provided checkpoints have many highly similar motions and face distortions as ground truth at the same time steps during testing, indicating possible training on the test set.
@@ 4.4 Quantitative Results
Impact of FOPPAS and Runtime Analysis. Diffusion models suffer from the long sampling steps at the test time – the original DDPM model takes 1000 steps to sample a quality image. Considering the resampling of Repaint [29], it would be at least 2 times slower. However, thanks to our FOPPAS mechanism, we can synthesize arbitrary-long smooth motion sequences in real-time. We test the runtime of our method on a one-minute-long audio corresponding to 900 frames of expressions and gestures (BEAT uses 15 FPS motion), with the number of overlapping frames in FOPPAS set to 4. The average runtime of our method is 28.6 seconds on a single NVIDIA GeForce RTX 3090 GPU, corresponding to around 31.5 FPS. The time for audio encoding (Mel-Spectrogram and HuBERT) is also included. In contrast, if we directly use DDPM with Repaint [29], it costs 2068.1s to infer 900 frames using the default setting of repaint, which is 0.44 FPS. Therefore, our proposed FOPPAS can benefit the diffusion model to achieve real-time in the inference stage, making many downstream products applicable.
@@ 4.5 Ablation Study
Uni-directional condition flow. To demonstrate the effectiveness of our uni-directional condition flow design, we conduct experiments on three ablated architectures: ours without expression condition flow into gesture encoder (Ours w/o $\mathbf{\hat{x}_{0}^{E}}$), ours without cutting off the gradient when feeding predicted expression to gesture encoder (Ours w/o Detach), naïvely concatenate expression and gesture together to be fed into a single Transformer denoising block (Ours Naïve Concat), and reverse the direction of condition flow to become gesture to expression (Ours Reverse Direction). For Ours Naïve Concat, we enlarge the latent dimension of the Transformer block by 1.41 times to have a similar computational complexity for a fair comparison. From Table 1, on both BEAT and SHOW datasets, our DiffSHEG outperforms all the four ablated architectures on Fréchet distances and achieves better or comparable BA, Diversity, and PCM/SRGR scores. On the BEAT dataset, the reverse direction ablation has a higher diversity due to the high entropy of gestures as a condition for expression, but its other metrics are still worse than DiffSHEG, indicating the harmful effect of gesture condition flow to expression. These results demonstrate the effectiveness of our design and validate our intuitive analysis of the relationship between expression and gesture.
Qualitative ablation. As shown in Figure 9, compared to other baselines, our full version of DiffSHEG usually exhibits more diversified and meaningful gestures, including head motions. For example, when saying “… for you”, our full version of DiffSHEG generates the motion with a large stress gesture, while other ablations only show a small motion. When saying “there are internal visions …”, our full version of motion shows a overturn and stop of hand as well as a turn of head, which corresponds to the speech stress on the word “internal”. In contrast, other ablations have hands remain still or relatively small motions.
@@ 6 Acknowledgement
We sincerely thank the help from Shaojun Liu and Zhaoxin Fan on the Metahuman character rendering in Unreal Engine.
@@ Appendix A User Study
We utilize the Tencent Questionnaire system to conduct the user study. Figure 10 is a screenshot of part of our questionnaire. The users are asked to download the videos to their local disk to avoid potential latency and out-of-synchronization between audio and video. The video order is shuffled to avoid inertia bias. Figure 11 is a screenshot of our shuffled video. The average time to complete the questionnaire is 40 min. To avoid noisy data, We dropped the questionnaires that took shorter than 10 min.
@@ Appendix B Algorithm for FOPPAS
To clearly describe the process of FOPPAS, we provide Algorithm 1 for reference. Algorithm 1 describes the single clip pass of our FOPPAS method in Sec. 3.5. Specifically, we utilize Repaint [29] as our out-painting method and adapt it to DDIM sampling to boost the inference speed of our method.
@@
Figure 10: Part of the questionnaire of our user study. We ask users to sort the three results according to the metric.
@@ Appendix D Analysis on Agility and Smoothness
We find our method shows the same level of agility and smoothness as the real motions. In contrast, most baselines fail to match the realistic level of agility and smoothness. In addition to watching the motion video to check the agility and smoothness, we also give a simple visualization of numerical values of motion, velocity, and acceleration. As shown in Figure 12(b) and Figure 12(c), we can obviously see that the velocity and acceleration of CaMN and DiffuseStyleGesture are consistently lower than GT, indicating lower agility. The DiffGesture has severe jittering gestures as shown in 12(a), therefore we did not visualize it in velocity and acceleration figures. Our method and ListenDenoiseAction have a comparative velocity as ground truth (GT) gestures, however, ListenDenoiseAction sometimes shows unstable motions such as at frame index of 195, 210, 250, and 260. As for smoothness, we can see CaMN has a very smooth curve in the motion Figure 12(a), however, it is too smooth compared to real motions. DiffuseStyleGesture and ListenDenoiseAction have comparative motion curve smoothness with ours and real motion, but if watch the motion video, we can still perceive that our method has the best smoothness compared to baselines, which is consistent with the results of user study.
@@ Appendix F Additional Implementation Details
Appendix G Limitations

### 06 Apr 2024 v2

@@ 4.3 Qualitative Results
Qualitative comparisons. We highly recommend readers watch our demo video to get an intuitive sense of qualitative results. Overall, in addition to better speech-motion alignment, we observe more realistic, agile and diverse expressions and gestures of ours than that of baselines on both datasets (Figure 456). On the BEAT dataset, our method shows more agile gestures than CaMN. The gestures of baseline diffusion-based methods all suffer different extents of jittering. The jittering extents from high to low are DiffGesture, DiffuseStyleGesture, and LDA. Both our method and CaMN can generate smooth motions. However, CaMN exhibits much slower and less diverse motion while ours is as agile as the real motion. For expression, CaMN shows inaccurate lip sync while other methods all look fine. Interestingly, we find that the CaMN barely blinks, while our method can generate good expressions with a reasonable frequency of blinks. Other baselines show a more frequent blinking, corresponding to higher diversity scores. However, ours has the most similar blink frequency as the real data. On the SHOW dataset, our method shows more agile gestures than TalkSHOW and more diverse expressions and gestures (Figure 56). This might be due to the fact that Talkshow utilizes VQ-VAE to encode the gestures, which may filter the high-frequency gestures, leading to slow motion. Also, because VQ-VAE tokenizes the gestures into limited numbers, its gestures show lower diversity than ours. The LS3DCG suffers from jittering arms, which may be because their deterministic CNN structure has difficulty approximating the many-to-many mapping. Note that we retrain the TalkSHOW for qualitative comparison because the results of TalkSHOW and LS3DCG with author-provided checkpoints have many highly similar concurrent motions and face distortions as ground truth during testing.
@@ 4.4 Quantitative Results
Impact of FOPPAS and Runtime Analysis. Diffusion models suffer from the long sampling steps at the test time – the original DDPM model takes 1000 steps to sample a quality image. Considering the resampling of Repaint [30], it would be slower by times. However, thanks to our FOPPAS mechanism, we can synthesize arbitrary-long smooth motion sequences in real time. We test the runtime of our method on a one-minute-long audio corresponding to 900 frames of expressions and gestures (BEAT uses 15 FPS motion), with the number of overlapping frames in FOPPAS set to 4. The average runtime of our method is 28.6 seconds on a single NVIDIA GeForce RTX 3090 GPU, corresponding to around 31.5 FPS. The time for audio encoding (Mel-Spectrogram and HuBERT) is also included. In contrast, if we directly use DDPM with Repaint [30], it costs 2068.1s to infer 900 frames using the default setting of repaint, which is 0.44 FPS. Therefore, our proposed FOPPAS can benefit the diffusion model to achieve real-time in the inference stage, making many downstream products applicable.
@@ 4.5 Ablation Study
Uni-directional condition flow. To demonstrate the effectiveness of our uni-directional condition flow design, we conduct experiments on three ablated architectures: ours without expression condition flow into gesture encoder (Ours w/o $\mathbf{\hat{x}_{0}^{E}}$), ours without cutting off the gradient when feeding predicted expression to gesture encoder (Ours w/o Detach), naïvely concatenate expression and gesture together to be fed into a single Transformer denoising block (Ours Naïve Concat), and reverse the direction of condition flow to become gesture to expression (Ours Reverse Direction). For Ours Naïve Concat, we enlarge the latent dimension of the Transformer block by 1.41 times to have a similar computational complexity for a fair comparison. From Table 1, on both BEAT and SHOW datasets, our DiffSHEG outperforms all the four ablated architectures on Fréchet distances and achieves better or comparable BA, Diversity, and PCM/SRGR scores. On the BEAT dataset, the reverse direction ablation has a higher diversity due to the high entropy of gestures as a condition for expression, but its other metrics are still worse than DiffSHEG, indicating the adverse affect of gesture condition flow to expression. These results demonstrate the effectiveness of our design and validate our intuitive analysis of the relationship between expression and gesture. For more ablations, please refer to our Appendix.
@@ 6 Acknowledgement
We sincerely thank the help from Shaojun Liu and Zhaoxin Fan on the MetaHuman character rendering.
@@ Appendix A User Study
We utilize the Tencent Questionnaire system to conduct the user study. Figure 8 is a screenshot of part of our questionnaire. The users are asked to download the videos to their local disk to avoid potential latency and out-of-synchronization between audio and video. The video order is shuffled to avoid inertia bias. Figure 9 is a screenshot of our shuffled video. The average time to complete the questionnaire is 40 min. To avoid invalid data, We dropped the questionnaires that took shorter than 10 min.
@@ Appendix B Algorithm for FOPPAS
To clearly describe the process of FOPPAS, we provide Algorithm 1 for reference. Algorithm 1 describes the single clip pass of our FOPPAS method in Sec. 3.5. Specifically, we utilize Repaint [30] as our out-painting method and adapt it to DDIM sampling to boost the inference speed.
@@
Figure 8: Part of the questionnaire of our user study. We ask users to sort the results of different methods according to the metrics.
@@ Appendix D Analysis on Agility and Smoothness
We find our method shows the same level of agility and smoothness as the real motions. In contrast, most baselines fail to match the realistic level of agility and smoothness.
Metrics for agility and smoothness. As shown in Table 2, we compute three metrics for smoothness and agility: AE, Vel, and MLVS, which are the mean acceleration error, mean velocity, and mean local velocity standard deviation of 3D joint positions, respectively. Our DiffSHEG and CaMN have the smallest AE, corresponding to the best smoothness. On the other hand, our DiffSHEG has the closest Vel and MLVS to GT, while CaMN and DSG show very low Vel and MLVS, which is consistent with the observation of their slow motions. Similarly, the jitter motion of LDA and DG leads to high values in three metrics.
Numerical value visualization. In addition to watching the motion video to check the agility and smoothness, we also give a simple visualization of numerical values of motion, velocity, and acceleration. As shown in Figure 10(b) and Figure 10(c), we can obviously see that the velocity and acceleration of CaMN and DiffuseStyleGesture are consistently lower than GT, indicating lower agility. The DiffGesture has severe jittering gestures as shown in 10(a), therefore we did not visualize it in velocity and acceleration figures. Our method and ListenDenoiseAction have a comparative velocity as ground truth (GT) gestures, however, ListenDenoiseAction sometimes shows unstable motions such as at frame index of 195, 210, 250, and 260. As for smoothness, we can see CaMN has a very smooth curve in the motion Figure 10(a), however, it is too smooth compared to real motions. DiffuseStyleGesture and ListenDenoiseAction have comparative motion curve smoothness with ours and real motion, but if watch the motion video, we can still perceive that our method has the best smoothness compared to baselines, which is consistent with the results of user study.
GT Ours CaMN DSG LDA DG
AE $\downarrow$ 0 0.628 0.617 0.637 1.213 5.210
Vel $\dagger$ 0.863 0.821 0.407 0.530 1.2531 3.399
[…]
Table 2: Smoothness and agility metric results. $\downarrow$: smaller is better. $\dagger$: closer to GT is better.
@@ Appendix E Metrics Computation
Motion Expression Gesture
[…]
Ours 324.67 331.72 0.1185 0.9400 438.93 0.9139 0.2507
w/o Hubert 347.07 325.91 0.1612 1.1079 495.86 0.8916 0.2228
w/o Mel 393.47 396.60 0.1297 0.9935 485.10 0.8933 0.2071
w/o Mid 437.23 450.85 0.1193 0.9464 613.86 0.8933 0.2210
[…]
Table 3: Audio encoder ablation results. $\downarrow$: smaller is better. $\dagger$: closer to GT is better.
Ours CaMN DG DSG LDA
[…]
L1$\downarrow$ 0.9400 1.2931 1.2002 1.3077 1.1383
Table 4: Lipsync results. $\downarrow$: smaller is better.
@@ Appendix F Additional Implementation Details
Appendix G Additional Experiments
Qualitative ablation. As shown in Figure 12, compared to other baselines, our full version of DiffSHEG usually exhibits more diversified and meaningful gestures, including head motions. For example, when saying “… for you", our full version of DiffSHEG generates the motion with a large stress gesture, while other ablations only show a small motion. When saying “there are internal visions …", our full version of motion shows a overturn and stop of hand as well as a turn of head, which corresponds to the speech stress on the word “internal". In contrast, other ablations have hands remain still or relatively small motions.
Audio encoder ablation. As shown in Table 3, our final version gets overall better results with three ablations: removing the Hubert feature (w/o Hubert), removing Mel Spectrogram (w/o Mel), and removing the trainable audio Transformer which learns the mid-level features (w/o Mid). The results indicate that removing any one of the three components will lead to a performance drop, especially the learnable mid-level audio transformer encoder. In addition, it seems that the low-level audio feature (Mel) has more influence on expression and the high-level audio feature (Hubert) affects more on gesture metrics.
Lip synchronous metric. We compute the lip-sync error in Table 4. According to the results, our DiffSHEG achieves better lip-sync scores than all the baselines.
Appendix H Limitations